OpenStudy (anonymous):

type and ordered pair used method of substitution to solve the system of linear equation 2x-y=6 and 7x-3y=22 i need help bad and I really don't what i'm doing so please help me

OpenStudy (anonymous):

you should take the value of 1 variable and put it in other like from 1st equation you should take the value 2x-y=6 . 2x=y+6 x=(y+6)/2 and put it in 2nd equation then solve it for y you will get the value of y then substitute it in any f the 2 equations and then find the value of x

hero (hero):

7x-3y=22 2x-y=6 Multiply both sides of 2x - y = 6 by three to get: 7x - 3y = 22 6x - 3y = 18 Subtract the second equation from the first to get: x = 4 Substitute x back into the original equation to find y

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

CAN U PLZ HELP ME